# Complexity of deep computations via topology of function spaces

### v1

@@ 3.1. NIP, Rosenthal compacta, and deep computations
Let $(L,\mathcal{P},\Gamma)$ be a compositional computational structure (Definition 2.3) satisfying the Extendability Axiom (Definition 2.4) with $\mathcal{P}$ countable. Let $R$ be an exhaustive collection of sizers. Let $\Delta\subseteq\Gamma$ be $R$-confined. The following are equivalent.
@@
Since $\mathcal{P}$ is countable, $\mathcal{L}[r_{\bullet}]\subseteq\mathbb{R}^{\mathcal{P}}$ is Polish. Also, the Extendability Axiom implies that $\pi_{P}\circ\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}$ is a pointwise bounded set of continuous functions for all $P\in\mathcal{P}$. Hence, Theorem 1.12 and Lemma 1.13 prove the equivalence of (i) and (ii). If (i) holds and $f\in\overline{\Delta}$, then write $f=\mathcal{U}{\rm lim}_{i}\gamma_{i}$ as an ultralimit. Define $\tilde{f}:=\mathcal{U}{\rm lim}_{i}\tilde{\gamma}_{i}$. Hence, for all $r_{\bullet}\in R$ we have $\tilde{f}|_{\mathcal{L}[r_{\bullet}]}\in\overline{\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}}\subseteq B_{1}(\mathcal{L}[r_{\bullet}],\mathcal{L}[r_{\bullet}])$. That every deep computation is a pointwise limit of a countable sequence of computations follows from the fact that for a Polish space $X$ every compact subset of $B_{1}(X)$ is Fréchet-Urysohn (that is, a space where topological closures coincide with sequential closures, see Theorem 3F in [BFT_1978_PCompactBaire] or Theorem 4.1 in [debs2013rosenthal]). ∎
@@ 3.2. The Todorčević trichotomy and levels of PAC learnability
Given a countable set $\Delta$ of computations satisfying the NIP on features and shards (condition (ii) of Theorem 3.1) we have that $\overline{\tilde{\Delta}_{\mathcal{L}[r_{\bullet}]}}$ (for a fixed sizer $r_{\bullet}$) is a separable Rosenthal compactum (see Definition 1.9). Todorčević proved a remarkable trichotomy for Rosenthal compacta [Todorcevic_1999_CompactSubsetsBaire], and later Argyros, Dodos, Kanellopoulos [argyros2008rosenthal] proved an heptachotomy that refined Todorčević’s classification. In this section, inspired by the work of Glasner and Megrelishvili [glasner2022tame], we study ways in which this classification allows us to obtain different levels of PAC-learnability and NIP.
Recall that a topological space $X$ is hereditarily separable if every subspace is separable, and that $X$ is first countable if every point in $X$ has a countable local basis. Every separable metrizable space is hereditarily separable, and R. Pol proved that every hereditarily separable Rosenthal compactum is first countable (see section 10 of [debs2013rosenthal]). This suggests the following definition:
@@
Let $(L,\mathcal{P},\Gamma)$ be a CCS satisfying the Extendability Axiom and $R$ be an exhaustive collection of sizers. Let $\Delta\subseteq\Gamma$ be an $R$-confined countable set of computations satisfying the NIP on shards and features (condition (ii) in Theorem 3.1). We say that $\Delta$ is:
@@
We now present some separable and non-separable examples of Rosenthal compacta.
@@
    Extended Alexandroff compactification. For each finite binary sequence $s\in 2^{<\mathbb{N}}$, let $v_{s}:2^{\mathbb{N}}\rightarrow\mathbb{R}$ be given by $v_{s}(x)=1$ if $x$ extends $s$ and $v_{s}(x)=0$ otherwise. Let $\hat{A}(2^{\mathbb{N}})$ be the pointwise closure of $\{v_{s}:s\in 2^{<\mathbb{N}}\}$, i.e., $\hat{A}(2^{\mathbb{N}})=A(2^{\mathbb{N}})\cup\{v_{s}:s\in 2^{<\mathbb{N}}\}$. Note that this space is a separable Rosenthal compactum which is not first countable.
@@
    Split Cantor. Let << be the lexicographic order in the space of infinite binary sequences, i.e., $2^{\mathbb{N}}$. For each $a\in 2^{\mathbb{N}}$ let $f_{a}^{-}:2^{\mathbb{N}}\rightarrow\mathbb{R}$ be given by $f_{a}^{-}(x)=1$ if $x<a$ and $f_{a}^{-}(x)=0$ otherwise. Let $f_{a}^{+}:2^{\mathbb{N}}\rightarrow\mathbb{R}$ be given by $f_{a}^{+}(x)=1$ if $x\leq a$ and $f_{a}^{+}(x)=0$ otherwise. The split Cantor is the space $S(2^{\mathbb{N}})=\{f_{a}^{-}:a\in 2^{\mathbb{N}}\}\cup\{f_{a}^{+}:a\in 2^{\mathbb{N}}\}$. This is a separable Rosenthal compactum. One example of a countable dense subset is the set of all $f_{a}^{+}$ and $f_{a}^{-}$ where $a$ is an infinite binary sequence that is eventually constant. Moreover, it is hereditarily separable, but it is not metrizable.
@@
    Alexandroff Duplicate. Let $K$ be any compact metric space and consider the Polish space $X=C(K)\sqcup K$, i.e., the disjoint union of $C(K)$ (with its supremum norm topology) and $K$. For each $a\in K$ define $g_{a}^{0},g_{a}^{1}:X\rightarrow\mathbb{R}$ as follows: g_a^0(x)={x(a),x∈C(K)0,x∈K g_a^1(x)={x(a),x∈C(K)δ_a(x),x∈K Let $D(K)=\{g_{a}^{0}:a\in K\}\cup\{g_{a}^{1}:a\in K\}$. Notice that $D(K)$ is a first countable Rosenthal compactum. It is not separable if $K$ is uncountable. The interesting case will be when $K=2^{\mathbb{N}}$.
@@
    Extended Alexandroff Duplicate of the split Cantor. For each finite binary sequence $t\in 2^{<\mathbb{N}}$ let $a_{t}\in 2^{\mathbb{N}}$ be the sequence starting with $t$ and ending with 0’s and let $b_{t}\in 2^{\mathbb{N}}$ be the sequence starting with $t$ and ending with $1$’s. Define $h_{t}:2^{\mathbb{N}}\rightarrow\mathbb{R}$ by h_t(x)={0,x¡a_t1/2,a_t≤x ≤b_t1,b_t¡x Let $\hat{D}(S(2^{\mathbb{N}}))$ be the pointwise closure of the set $\{h_{t}:t\in 2^{<\mathbb{N}}\}$. Hence, $\hat{D}(S(2^{\mathbb{N}}))$ is a separable first countable Rosenthal compactum which is not hereditarily separable. In fact, it contains an uncountable discrete subspace (see Theorem 5 in [Todorcevic_1999_CompactSubsetsBaire]).
Theorem 3.4 (Todorčević’s Trichotomy, [Todorcevic_1999_CompactSubsetsBaire], Theorem 3 in [argyros2008rosenthal]).
@@
In the three separable three cases given in 3.3, namely, ($\hat{A}(2^{\mathbb{N}})$, $S(2^{\mathbb{N}})$ and $\hat{D}(S(2^{\mathbb{N}}))$), the countable dense subsets are indexed by the binary tree $2^{<\mathbb{N}}$. This choice of index is useful for two reasons:
@@
    Our emphasis is computational. Elements of $2^{<\mathbb{N}}$ represent finite bitstrings, i.e., standard computations, while Rosenthal compacta represent deep computations, i.e., limits of finite computations. Mathematically, deep computations are pointwise limits of standard computations. Howwever, computationally, we are interested in the manner (and the efficiency) in which the approximations can occur.
@@
    The Ramsey theory of perfect subsets of the Cantor space $2^{\mathbb{N}}$ can be imported to analyze the behavior of the accumulation points. Since $2^{<\mathbb{N}}$ is countable, we can always choose this index for the countable dense subsets. This is done in [argyros2008rosenthal].
@@
One of the main results in [argyros2008rosenthal] is that, up to equivalence, there are seven minimal families of Rosenthal compacta and that for every relatively compact $\{f_{t}:t\in 2^{<\mathbb{N}}\}\subseteq B_{1}(X)$ there is a dyadic subtree $\{s_{t}:t\in 2^{<\mathbb{N}}\}$ such that $\{f_{s_{t}}:t\in 2^{<\mathbb{N}}\}$ is equivalent to one of the minimal families. We shall describe the seven minimal families next. We follow the same notation as in [argyros2008rosenthal]. For any node $t\in 2^{<\mathbb{N}}$, let us denote by $t^{\frown}0^{\infty}$ ($t^{\frown}1^{\infty}$) the infinite binary sequence starting with $t$ and continuing will all 0’s (respectively, all $1$’s). Fix a regular dyadic subtree $R=\{s_{t}:t\in 2^{<\mathbb{N}}\}$ of $2^{<\mathbb{N}}$ (i.e., a dyadic subtree such that every level of $R$ is contained in a level of $2^{<\mathbb{N}}$) with the property that for all $s,s^{\prime}\in R$, $s^{\frown}0^{\infty}\neq s^{\prime\frown}0^{\infty}$ and $s^{\frown}1^{\infty}\neq s^{\prime\frown}1^{\infty}$. Given $t\in 2^{<\mathbb{N}}$, let $v_{t}$ be the characteristic function of the set $\{x\in 2^{\mathbb{N}}:x\text{ extends }t\}$. Let << be the lexicographic order in $2^{\mathbb{N}}$. Given $a\in 2^{\mathbb{N}}$, let $f^{+}_{a}:2^{\mathbb{N}}\rightarrow\{0,1\}$ be the characteristic function of $\{x\in 2^{\mathbb{N}}:a\leq x\}$ and let $f^{-}_{a}:2^{\mathbb{N}}\rightarrow\{0,1\}$ be the characteristic function of $\{x\in 2^{\mathbb{N}}:a<x\}$. Given two maps $f,g:2^{\mathbb{N}}\rightarrow\mathbb{R}$ we denote by $(f,g):2^{\mathbb{N}}\sqcup 2^{\mathbb{N}}\rightarrow\mathbb{R}$ the function which is $f$ on the first copy of $2^{\mathbb{N}}$ and $g$ on the second copy of $2^{\mathbb{N}}$.
@@
Theorem 3.7 (Heptachotomy of minimal families, Theorem 2 in [argyros2008rosenthal]).
@@
4. Measure-theoretic versions of NIP and universal Monte Carlo computability of deep computations
The countability assumption on $\mathcal{P}$ played a crucial role in the proof of Theorem 1.12 , as it makes $\mathbb{R}^{\mathcal{P}}$ a Polish space. In this section, we do not assume that $\mathcal{P}$ is countable. We replace deterministic computability by measure-theoretic (‘Monte Carlo’) computability.
4.1. A measure-theoretic version of NIP
Recall that the raison d’être of the class of Baire-1 functions is to have a class that contains the continuous functions but is closed under pointwise limits, and that for perfectly normal $X$, a function $f$ is in $B_{1}(X,Y)$ if and only if $f^{-1}[U]$ is an $F_{\sigma}$ subset of $X$ for every open $U\subseteq Y$ (see Fact 1.2). This motivates the following definition:

### v2

@@ 3.1. NIP, Rosenthal compacta, and deep computations
Let $(L,\mathcal{P},\Gamma)$ be a compositional computational structure (Definition 2.3) satisfying the Extendibility Axiom (Definition 2.5) with $\mathcal{P}$ countable. Let $R$ be an exhaustive collection of sizers. Let $\Delta\subseteq\Gamma$ be $R$-confined. The following are equivalent.
@@
Since $\mathcal{P}$ is countable, $\mathcal{L}[r_{\bullet}]\subseteq\mathbb{R}^{\mathcal{P}}$ is Polish. Also, the Extendibility Axiom implies that $\pi_{P}\circ\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}$ is a pointwise bounded set of continuous functions for all $P\in\mathcal{P}$. Hence, Theorem 1.11 and Lemma 1.12 prove the equivalence of (i) and (ii). If (i) holds and $f\in\overline{\Delta}$, then write $f=\mathcal{U}{\rm lim}_{i}\gamma_{i}$ as an ultralimit. Define $\tilde{f}:=\mathcal{U}{\rm lim}_{i}\tilde{\gamma}_{i}$. Hence, for all $r_{\bullet}\in R$ we have $\tilde{f}|_{\mathcal{L}[r_{\bullet}]}\in\overline{\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}}\subseteq B_{1}(\mathcal{L}[r_{\bullet}],\mathcal{L}[r_{\bullet}])$. That every deep computation is a pointwise limit of a countable sequence of computations follows from the fact that for a Polish space $X$ every compact subset of $B_{1}(X)$ is Fréchet-Urysohn (that is, a space where topological closures coincide with sequential closures, see Theorem 3F in [BFT78] or Theorem 4.1 in [DEB13]). ∎
@@ 3.2. The Todorčević trichotomy and levels of PAC learnability
In this subsection we study the case when the set of deep computations is a separable Rosenthal compactum. We are interested in the separable case for two reasons:
  1. (1)
    In practice, the set $\Delta$ of computations is countable. This implies that the set $\overline{\Delta}$ of deep computations is separable.
  2. (2)
    The non-separable case lacks some tools and nice examples, which makes their study more complicated. In the separable case we have two important results, which are introduced in this subsection (Todorčević’s Trichotomy) and the next subsection (Argyros-Dodos-Kanellopoulos heptachotomy). By introducing Todorčević’s Trichotomy into this framework, we obtain a classification of the complexity of deep computations.
Given a countable set $\Delta$ of computations satisfying the NIP on features and shards (condition (ii) of Theorem 3.1) we have that $\overline{\tilde{\Delta}_{\mathcal{L}[r_{\bullet}]}}$ (for a fixed sizer $r_{\bullet}$) is a separable Rosenthal compactum (see Definition 1.9). Todorčević proved a remarkable trichotomy for Rosenthal compacta [TOD99], and later Argyros, Dodos, Kanellopoulos [ADK08] proved an heptachotomy that refined Todorčević’s classification. In this section, inspired by the work of Glasner and Megrelishvili [GM22], we study ways in which this classification allows us to obtain different levels of PAC-learnability and NIP.
Recall that a topological space $X$ is hereditarily separable if every subspace is separable, and that $X$ is first countable if every point in $X$ has a countable local basis. Every separable metrizable space is hereditarily separable, and R. Pol proved that every hereditarily separable Rosenthal compactum is first countable (see section 10 of [DEB13]). This suggests the following definition:
@@
Let $(L,\mathcal{P},\Gamma)$ be a CCS satisfying the Extendibility Axiom and $R$ be an exhaustive collection of sizers. Let $\Delta\subseteq\Gamma$ be an $R$-confined countable set of computations satisfying the NIP on shards and features (condition (ii) in Theorem 3.1). We say that $\Delta$ is:
@@
We now present some separable and non-separable examples of Rosenthal compacta. These show that the previously discussed classes NIPi are not equal.
@@
    Extended Alexandroff compactification. For each finite binary sequence $s\in 2^{<\mathbb{N}}$, let $v_{s}:2^{\mathbb{N}}\rightarrow\mathbb{R}$ be given by $v_{s}(x)=1$ if $x$ extends $s$ and $v_{s}(x)=0$ otherwise. Let $\hat{A}(2^{\mathbb{N}})$ be the pointwise closure of $\{v_{s}:s\in 2^{<\mathbb{N}}\}$, i.e., $\hat{A}(2^{\mathbb{N}})=A(2^{\mathbb{N}})\cup\{v_{s}:s\in 2^{<\mathbb{N}}\}$. Note that this space is a separable Rosenthal compactum which is not first countable. This is the example discussed in Section 2.5. It is an example of a CCS that is NIP but not NIP1.
@@
    Split Cantor. Let << be the lexicographic order in the space of infinite binary sequences, i.e., $2^{\mathbb{N}}$. For each $a\in 2^{\mathbb{N}}$ let $f_{a}^{-}:2^{\mathbb{N}}\rightarrow\mathbb{R}$ be given by $f_{a}^{-}(x)=1$ if $x<a$ and $f_{a}^{-}(x)=0$ otherwise. Let $f_{a}^{+}:2^{\mathbb{N}}\rightarrow\mathbb{R}$ be given by $f_{a}^{+}(x)=1$ if $x\leq a$ and $f_{a}^{+}(x)=0$ otherwise. The split Cantor is the space $S(2^{\mathbb{N}})=\{f_{a}^{-}:a\in 2^{\mathbb{N}}\}\cup\{f_{a}^{+}:a\in 2^{\mathbb{N}}\}$, which was obtained as the closure of the space discussed in Section 2.4, giving an example separating NIP2 from NIP3. This is a separable Rosenthal compactum. One example of a countable dense subset is the set of all $f_{a}^{+}$ and $f_{a}^{-}$ where $a$ is an infinite binary sequence that is eventually constant. Moreover, it is hereditarily separable, but it is not metrizable.
@@
    Alexandroff Duplicate. Let $K$ be any compact metric space and consider the Polish space $X=C(K)\sqcup K$, i.e., the disjoint union of $C(K)$ (with its supremum norm topology) and $K$. For each $a\in K$ define $g_{a}^{0},g_{a}^{1}:X\rightarrow\mathbb{R}$ as follows: g_a^0(x)={x(a), if x∈C(K);0, if x∈K; g_a^1(x)={x(a), if x∈C(K);δa(x), if x∈K. Let $D(K)=\{g_{a}^{0}:a\in K\}\cup\{g_{a}^{1}:a\in K\}$. Notice that $D(K)$ is a first countable Rosenthal compactum. It is not separable if $K$ is uncountable. The interesting case will be when $K=2^{\mathbb{N}}$.
@@
    Extended Alexandroff Duplicate of the split Cantor. For each finite binary sequence $t\in 2^{<\mathbb{N}}$ let $a_{t}\in 2^{\mathbb{N}}$ be the sequence starting with $t$ and ending with 0’s and let $b_{t}\in 2^{\mathbb{N}}$ be the sequence starting with $t$ and ending with $1$’s. Define $h_{t}:2^{\mathbb{N}}\rightarrow\mathbb{R}$ by h_t(x)={0, if x¡at;1/2, if at≤x ≤bt;1, if bt¡x. Let $\hat{D}(S(2^{\mathbb{N}}))$ be the pointwise closure of the set $\{h_{t}:t\in 2^{<\mathbb{N}}\}$. Hence, $\hat{D}(S(2^{\mathbb{N}}))$ is a separable first countable Rosenthal compactum which is not hereditarily separable. In fact, it contains an uncountable discrete subspace (see Theorem 5 in [TOD99]).
Theorem 3.4 (Todorčević’s Trichotomy, [TOD99], Theorem 3 in [ADK08]).
@@
Todorčević’s Trichotomy suggests that in order to distinguish the classes NIPi, the examples in 3.3 are essential. The following examples show that the levels NIPi ($i=1,2,3$) may be distinguished by the topological complexity of deep computations.
Examples 3.5.
  1. (1)
    Let $(L,\mathcal{P},\Gamma)$ be the computation of square root (example 2.6 with $\Delta=\Gamma$. We saw that $\overline{\tilde{\Delta}}=\tilde{\Delta}\cup\{\tilde{f}\}\subseteq B_{1}(\mathcal{L},\mathcal{L})$. This corresponds to the Alexandroff compactification of a countable discrete set, which is metrizable. Hence, $\Delta$ is NIP3 but it is not stable, in the sense that $\overline{\tilde{\Delta}}\not\subseteq C(\mathcal{L},\mathcal{L})$.
  2. (2)
    Let $(L,\mathcal{P},\Gamma)$ be the finite precision threshold classifiers (Section 2.4) with $\Delta=\{\phi_{w}:w\in 2^{<\mathbb{N}}\}\cup\{\mathbf{0^{\infty}},\mathbf{1^{\infty}}\}$. We saw that $\overline{\Delta}$ is homeomorphic to the Split Cantor space $S(2^{\mathbb{N}})$ (Example 3.3(3)), which is hereditarily separable but not metrizable. Hence, $\Delta$ is NIP2 but not NIP3.
  3. (3)
    Let $(L,\mathcal{P},\Gamma)$ be the finite precision prefix test (Section 2.5) with $\Delta=\{\psi_{w}:w\in 2^{<\mathbb{N}}\}$. We saw that $\overline{\tilde{\Delta}}$ is homeomorphic to the Extended Alexandroff compactification $\hat{A}(2^{\mathbb{N}})$ (Example 3.3(3)), which is separable but not first countable. Hence, $\Delta$ is NIP but not NIP1.
@@
In the three separable cases given in 3.3, namely, $\hat{A}(2^{\mathbb{N}})$, $S(2^{\mathbb{N}})$ and $\hat{D}(S(2^{\mathbb{N}}))$, the countable dense subsets are indexed by the binary tree $2^{<\mathbb{N}}$. This choice of index is useful for two reasons:
@@
    Our emphasis is computational. Elements of $2^{<\mathbb{N}}$ represent finite bitstrings, i.e., standard computations, while Rosenthal compacta represent deep computations, i.e., limits of finite computations. Mathematically, deep computations are pointwise limits of standard computations. However, computationally, we are interested in the manner (and the efficiency) in which the approximations can occur.
@@
    Infinite branches of the binary tree $2^{<\mathbb{N}}$ correspond to the Cantor space $2^{\mathbb{N}}$, the canonical perfect set (in the sense that any Polish space with no isolated points contains a copy of $2^{\mathbb{N}}$). The use of infinite dimensional Ramsey theory for trees (pioneered by the work of James D. Halpern, Hans Läuchli in [HL66] and also Keith Milliken in [MIL81], and Alain Louveau, Saharon Shelah, Boban Velickovic in [LSV93]) and perfect sets (Fred Galvin and Andreas Blass in [BLA81], Arnold W. Miller in [MIL89], and Stevo Todorčević in [TOD99]) allowed S.A. Argyros, P. Dodos and V. Kanellopoulos in [ADK08] to obtain an improved version of Theorem 3.4. It is no surprise that Ramsey Theory becomes relevant in the study of Rosenthal compacta as it was a key ingredient in Rosenthal’s $\ell_{1}$ Theorem. For this reason, the main results in [ADK08] (which we cite in this paper) are better explained by indexing Rosenthal compacta with the binary tree.
@@
One of the main results in [ADK08] is that, up to equivalence, there are seven minimal families of Rosenthal compacta and that for every relatively compact $\{f_{t}:t\in 2^{<\mathbb{N}}\}\subseteq B_{1}(X)$ there is a dyadic subtree $\{s_{t}:t\in 2^{<\mathbb{N}}\}$ such that $\{f_{s_{t}}:t\in 2^{<\mathbb{N}}\}$ is equivalent to one of the minimal families. We shall describe the seven minimal families next. We follow the same notation as in [ADK08]. For any node $t\in 2^{<\mathbb{N}}$, let us denote by $t^{\frown}0^{\infty}$ ($t^{\frown}1^{\infty}$) the infinite binary sequence starting with $t$ and continuing will all 0’s (respectively, all $1$’s). Fix a regular dyadic subtree $R=\{s_{t}:t\in 2^{<\mathbb{N}}\}$ of $2^{<\mathbb{N}}$ (i.e., a dyadic subtree such that every level of $R$ is contained in a level of $2^{<\mathbb{N}}$) with the property that for all $s,s^{\prime}\in R$, $s^{\frown}0^{\infty}\neq s^{\prime\frown}0^{\infty}$ and $s^{\frown}1^{\infty}\neq s^{\prime\frown}1^{\infty}$. Given $t\in 2^{<\mathbb{N}}$, let $v_{t}$ be the characteristic function of the set $\{x\in 2^{\mathbb{N}}:x\text{ extends }t\}$. Let << be the lexicographic order in $2^{\mathbb{N}}$. Given $a\in 2^{\mathbb{N}}$, let $f^{+}_{a}:2^{\mathbb{N}}\rightarrow\{0,1\}$ be the characteristic function of $\{x\in 2^{\mathbb{N}}:a\leq x\}$ and let $f^{-}_{a}:2^{\mathbb{N}}\rightarrow\{0,1\}$ be the characteristic function of $\{x\in 2^{\mathbb{N}}:a<x\}$. Given two maps $f,g:2^{\mathbb{N}}\rightarrow\mathbb{R}$ we denote by $(f,g):2^{\mathbb{N}}\sqcup 2^{\mathbb{N}}\rightarrow\mathbb{R}$ the function which is $f$ on the first copy of $2^{\mathbb{N}}$ and $g$ on the second copy of $2^{\mathbb{N}}$.
@@
Theorem 3.8 (Heptachotomy of minimal families, Theorem 2 in [ADK08]).
@@
The implications of this result for deep computations is the following: for any countable set of computations $\Delta$ satisfying the NIP (for some CCS $(L,\mathcal{P},\Gamma)$), we can always find a countable discrete set of deep computations that approximate all the other deep computations. For example: in the finite precision prefix test example (subsection 2.5), the prefix test computations (family $D_{5}$) approximate all other deep computations. However, note that this discrete set $D_{i}$ may not consist of continuous functions (i.e., they will not be computable in general). For example, functions in $D_{3}$ are not continuous.
4. Randomized versions of NIP and Monte Carlo computability of deep computations
In this section, we replace deterministic computability by probabilistic (‘Monte Carlo’) computability. We do not assume that $\mathcal{P}$ is countable. The main results of the section are Theorem 4.8 (connecting NIP and Monte Carlo computability) and 4.14 (connecting Talagrand stability and Monte Carlo computability).
Fundamental in this section is a measure-theoretic version of Theorem 1.11, namely, Theorem 4.7. For the proof of Theorem 1.11, we assumed countability of $\mathcal{P}$ — this ensured that $\mathbb{R}^{\mathcal{P}}$ a Polish space. In this section, the countability assumption is not needed.
4.1. NIP and Monte Carlo computability of deep computations
The raison d’être of the Baire class-1 functions is have a class of functions that contains the continuous functions and is closed under pointwise limits. By Fact 1.2, for perfectly normal $X$, a function $f$ is in $B_{1}(X,Y)$ if and only if $f^{-1}[U]$ is an $F_{\sigma}$ subset of $X$ for every open $U\subseteq Y$. This motivates the following definition: